If anyone could help me with the following 2 problems I would appreciate it.
Thank You
Principal of zero
v^2+5v-14=0
-----------------------------------------------
The polynomial 0.041h-0.018a-2.69 can be used to estimate the lung capacity in liters of a female with a height in centimeters and age in years. Find the lung capacity of a 60 year old woman who is 170cm tall.
Her lung capacity is approximately _______ liters.

Answer by stanbon(75887)

Principal of zero
v^2+5v-14=0
(v-7)(v+2)=0
v = 7 or v=-2

-----------------------------------------------
The polynomial 0.041h-0.018a-2.69 can be used to estimate the lung capacity in liters of a female with a height in centimeters and age in years. Find the lung capacity of a 60 year old woman who is 170cm tall.
Her lung capacity is approximately _______ liters.
0.0141*170 - 0.0182*60 - 2.69 = -1.3850
---------
Comment: I think the polynomial has a typo somewhere ; the calculation
should not turn out to be negative.
=============================
Cheers,
Stan H.